5x +2y +z =-8


x -2y -3z =0


-x +y +2z =3


Solved this problem by using Gauess Gordan method.


1
Expert's answer
2021-11-30T13:37:06-0500

"\begin{cases}
5x+2y+z=-8
\\
x-2y-3z=0
\\
-x+y+2z=3
\end{cases}"


"\begin{bmatrix} 
5&2&1&|&-8
\\
1&-2&-3&|&0
\\
-1&1&2&|&3
\end{bmatrix}

\overset{R_3+R_2}{\rightarrow}
\begin{bmatrix} 
5&2&1&|&-8
\\
1&-2&-3&|&0
\\
0&-1&-1&|&3
\end{bmatrix}"


"\qquad \qquad \qquad \qquad \qquad \quad\overset{R_1-5R_2}{\rightarrow}

\begin{bmatrix} 
0&12&16&|&-8
\\
1&-2&-3&|&0
\\
0&-1&-1&|&3
\end{bmatrix}"


"\qquad \qquad \qquad \qquad \qquad \quad\overset{R_1+12R_3}{\rightarrow}

\begin{bmatrix} 
0&0&4&|&28
\\
1&-2&-3&|&0
\\
0&-1&-1&|&3
\end{bmatrix}"


"\qquad \qquad \qquad \qquad \qquad \quad\overset{R_2-2R_3}{\rightarrow}

\begin{bmatrix} 
0&0&4&|&28
\\
1&0&-1&|&-6
\\
0&-1&-1&|&3
\end{bmatrix}"


"\qquad \qquad \qquad \qquad \qquad \quad\overset{R_1/4}{\rightarrow}

\begin{bmatrix} 
0&0&1&|&7
\\
1&0&-1&|&-6
\\
0&-1&-1&|&3
\end{bmatrix}"


"\qquad \qquad \qquad \qquad \qquad \quad\overset{R_2+R_1}{\rightarrow}

\begin{bmatrix} 
0&0&1&|&7
\\
1&0&0&|&1
\\
0&-1&-1&|&3
\end{bmatrix}"


"\qquad \qquad \qquad \qquad \qquad \quad\overset{R_3+R_1}{\rightarrow}

\begin{bmatrix} 
0&0&1&|&7
\\
1&0&0&|&1
\\
0&-1&0&|&10
\end{bmatrix}"


"\qquad \qquad \qquad \qquad \qquad \quad\overset{-R_3}{\rightarrow}

\begin{bmatrix} 
0&0&1&|&7
\\
1&0&0&|&1
\\
0&1&0&|&-10
\end{bmatrix}"


Answer: "x=1,\ \ y=-10,\ \ z=7."